a) What’s the future value of $100 after 3 years if it earns 8%, annual compounding?

b) What’s the present value of $100 to be received in 3 years if the interest rate is 8%, annual compounding?

c) What annual interest rate would cause $1,000 to grow to $2,000 in 8 years?

d) If a company’s sales are growing at a rate of 7.2% annually, how long will it take sales to double?.

e) What is the present value of a 6-year, $100 ordinary annuity if the annual interest rate is 4%?
Answer #1

A. Present Value (PV) = $ 100 Time Period (N) = 3 Years Interest (R) = 8 % Future Value = PV *(1+R%)^N Future Value = 100 *(1

Annual Interest Rate R = ? Present Value (PV) = $1,000 Future Value (FV) = $2,000 Time Period (N) = 8 Years Sales to be doubl

E. To find present value of ordinary annuity Annual interest rate R=4% Time period = 6 Years Annity (Cash flow ) = $ 100 Pres